OpenStudy (anonymous):

Solve the following equation: 1/x+2/(x-1)=3

OpenStudy (jamesj):

Multiply both sides by x(x-1). After doing that, do you now see where you are?

OpenStudy (anonymous):

x = 1 x = 1/3

OpenStudy (anonymous):

I will go from here james j we shal see how i turn out

OpenStudy (anonymous):

haha i still don't know what to do

OpenStudy (jamesj):

Well ... x(x-1) ( 1/x + 2/(x-1) ) = x(x-1)/x + 2x(x-1)/(x-1) Agreed? Now what do you do ...

OpenStudy (anonymous):

i do not know

OpenStudy (anonymous):

genuinely

OpenStudy (jamesj):

Repost and see if someone else can help you. I need to step away.

OpenStudy (anonymous):

genuinely

OpenStudy (anonymous):

ok thank you